OpenStudy (anonymous):

a. Find the slope of the line passing through the points (7,–4) and (–6,–7).

OpenStudy (azureilai):

Slope is \[\frac{ \Delta y }{ \Delta x }\]

OpenStudy (anonymous):

woah. >3<

OpenStudy (azureilai):

Do you know what to do now?

OpenStudy (anonymous):

now madam.. :/

OpenStudy (alexwee123):

\[m=\frac{ y2-y1 }{ x2-x1}\] slope=m you have two points (x1,y1) and (x2,y2) input in the values \[m=\frac{ -7-(-4) }{ -6-7 }\]

OpenStudy (alexwee123):

\[m=\frac{ -3 }{ -13 }\] \[m=\frac{ 3 }{ ? }\]

OpenStudy (alexwee123):

m=3/13

OpenStudy (anonymous):

so the answe is 3/13?

OpenStudy (alexwee123):

yep got any questions?

OpenStudy (anonymous):

none.. but could you help me with this one? Solve the given linear system 2x-4y=-6 -x+2y=3

OpenStudy (azureilai):

you can use substitution for that one. Take one of the equations and isolte y to one side (im going to use second one) 2y=x+3 y=(x+3)/2 and now you substitue y in the first equation 2x-4((x=3)/2))=-6 Solve for x, and when you get x, plug it into the second equation and solve

OpenStudy (azureilai):

isolate*

OpenStudy (anonymous):

by the way if (x1,y1)& (x2,y2) is like this then the given is (7,–4) and (–6,–7).

OpenStudy (anonymous):

thanks...

OpenStudy (alexwee123):

yep^ and you can also try using elimination to solve it...

OpenStudy (anonymous):

how to do the elimination method?

OpenStudy (alexwee123):

you would add and subtract the equations to cancel out a variable to do this you may have to multiply an equation or two for them to cancel each other out but for this set of equations are the same so you'd have infinite solutions

OpenStudy (anonymous):

could you give me some sample?

OpenStudy (alexwee123):

ya check this site out http://www.mathportal.org/algebra/solving-system-of-linear-equations/elimination-method.php it gives you a pretty good step by step procedure

OpenStudy (anonymous):

thanks.. :D